OpenStudy (anonymous):

Let f(x) = 3x3 – 4x – 1 and g(x) = x + 1. Find f of x over g of x

OpenStudy (muhammad_nauman_umair):

f(x)/g(x)=3x^3 – 4x – 1/x+1

OpenStudy (anonymous):

can you show your work

OpenStudy (muhammad_nauman_umair):

According to the given question : f of x =f(x) & g of x =g(x) as f(x) = 3x^3 – 4x – 1 & g(x) = x + 1 then f of x over g of x means f(x)/g(x) so by dividing f(x) by g(x) we have the above mentioned result....